1. Calculate the employers payroll taxes for the week
2. Prepare the general journal entry to record the payroll taxes

Problem 3- Calculating and recording employer payroll taxes 18 points Shoreview, Inc. had the following for the week ended March 13, 20XX Gross earnings for the period FICA-OASDI taxable earnings FUTA and SUTA taxable earnings $146,750 $131,640 39,500 using the tax rates mentioned in the text, with a state unemployment tax rate of 2.9%. 1. Calculate the employers payroll taxes for the week 2. Prepare the general journal entry to record the payroll taxes
Answer #1

given data

gross earnings for the period = $ 146750

FICA taxable earnings = $ 131640

FUTA and SUTA tacable earnings = $ 39500

unemployment tax rate = 2.9%

1) computation of employee payroll taxes :

particulars amount
FICA taxable earnings 131640
FUTA taxable earnings 39500
FICA social decurity taxes (working note 1) 10070
federal unemployment taxes (FUTA) (working note 2) 316
state unemploymeny taxes (working note 3) 1146
total 182672

working note 1 :

FICA social security taxes = FICA taxable earnings * tax rate (standard rate)

= 131640 * 7.65%

= 10070

working note 2 :

federal unemployment taxes (FUTA) = FUTA taxable earning * tax rate (standard rate)

= 39500 * 0.8%

= 316

working note 3 :

state unemployment taxes (SUTA) = FUTA taxable earning * unemployment tax rate

= 39500 * 2.9%

= 1146

2) journal entry to record the payroll taxes :

date particulars l/f debit credit
--- payroll tax expense a/c dr 11532
to FICA taxes payable a/c 10070
to FUTA taxes payable a/c 316
to SUTA taxes payable a/c 1146
(being payroll tax expenses will be raised)